Use the given information to find the number of degrees of​ freedom, the critical values chi Subscript Upper L Superscript 2χ2L and chi Subscript Upper R Superscript 2χ2R​, and the confidence interval estimate of sigmaσ. It is reasonable to assume that a simple random sample has been selected from a population with a normal distribution. Platelet Counts of Women 99​% ​confidence; n=27​, s=65.9.
Answer #1

Solution :

Given that,

c = 0.99

s = 65.9

n = 27

At 99% confidence level the is ,

= 1 - 99% = 1 - 0.99 = 0.01

/ 2 = 0.01 / 2 = 0.005

/2,df = 0.005,26 = 48.29

and

1- /2,df = 0.995,26 = 11.16

2L = 2/2,df = 48.29

2R = 21 - /2,df = 11.16

The 99% confidence interval for is,

s (n-1) / /2,df < < s (n-1) / 1- /2,df

65.9( 27 - 1 ) / 48.29< < 65.9( 27- 1 ) / 11.16

48.36 < < 100.59

( 48.36 , 100.59)